can you please help me figure out the following problem??
solve using the addition method
-2x=4y-14
5y=31-4x
(__,__)

Found 2 solutions by scott8148, nospmoht:
Answer by scott8148(6628)

multiplying the first equation by 2 ___ -4x = 8y - 28

adding the second equation ___ -4x + 5y = 8y - 28 + 31 - 4x

adding 4x ___ 5y = 8y + 3 ___ y = -1

substituting ___ -2x = 4(-1) - 14 ___ x = 9

Answer by nospmoht(3)

First switch it around so that all the variables are on one side and the numbers are on the other:
2x+4y=14
4x+5y=31
Then you want to get one variable out of the equation (this is where the addition method comes nto play). Lets get rid of X since it's easiest to do. You have to multiply one of the equations so that the Xs cancel out and you are left with only a Y:
-2(2x+4y=14) -->> -4x-8y=-28
Add the equations: + 4x+5y= 31

3y= 3
Solve: y= 1
Put the answer to Y into one of the original equations and you'll find the answer:
-2x=4(1)-14
-2x=4-14
-2x=-10
x=5
Answer: (5,1)
Hope this helped you man. Enjoy :)